# The Nuclear Star Cluster of M 74: a fossil record of the very early stages of a star-forming galaxy

### v1

@@ 6.3 NSC formation and the host-galaxy evolution
Different NSC properties have been found in galaxies of different masses (Sect. 1). This was explained with a transition in the dominant mechanisms driving NSC formation and growth, with in-situ star formation being dominant in higher mass galaxies, and globular cluster migration to the center of a galaxy dominating at lower masses (e.g., Fahrion2021; Fahrion2022a; Fahrion2022b). Among others, one piece of evidence is the transition in the NSC metallicity regime at galaxy stellar masses of about $10^{9}$ M(e.g, Neumayer2020). In more massive galaxies, NSCs tend to be more metal-rich, with [Fe/H] metallicities about $\pm 0.5$ dex around solar. For lower galaxy masses, on the other hand, NSCs have metallicities in a wide range, including very low values. In Fig.7, we have reconstructed this trend as shown in Fig. 9 of Neumayer2020, and updated it with additional measurements (NSCs from Fahrion2021; Fahrion2022b; Kacharov2018). NSCs in LTGs are shown in blue, and show, on average, lower metallicities than NSCs in ETGs (red symbols). Note also that blue triangles are total metallicities ([M/H], from Fahrion2022b), and they have to be considered upper limits for [Fe/H] metallicities. While more points will be needed to shed more light on this trend, the available points suggest that the trend of more metal-rich NSCs being hosted by more massive galaxies, clearly shown for ETGs, also holds for LTGs (but with a lower average metallicity). For higher-mass spirals, we only have four NSCs from the literature (Kacharov2018, blue circles), which appear to display lower metallicities than, on average, NSCs in ETGs of similar masses.
@@
For M 74, [Fe/H] was calculated from [M/H] and [Mg/Fe] in Table 1 following the relation ${\rm[Fe/H]}={\rm[M/H]}-0.75{\rm[Mg/Fe]}$ from Vazdekis2015. M 74’s NSC (blue open circle), is more metal poor than the rest of NSCs in Fig. 7 for similar galaxy masses, even if compared with the few LTGs. M 74 is located, in Fig. 7, in a mass regime where NSCs are thought to be dominated by in-situ formation, as suggested by their usually higher metallicities than their host galaxies (Fahrion2022a). However, its host galaxy (blue open square) is much more metal-rich than the NSC (about 0.7 dex difference in [M/H] within the central PSF region). In general, NSC old ages and low metallicities, in particular if older and more metal-poor than the host galaxy, have been interpreted as a main formation through globular cluster inspiral (e.g., Fahrion2021). If compared with other star clusters in M 74, its NSC shows similar ages to the most massive globular clusters (Whitmore2023), supporting the cluster migration scenario. On the other hand, models predict that NSCs of the mass of M 74’s NSC (${\sim}10^{7}$ M) assembled slightly more than half of their mass via in-situ star formation and the rest through globular cluster inspiral (Fahrion2022a). Recent JWST studies at high redshift have recently proposed that similar very dense objects existed already in the early universe and may have been the first structure that a galaxy formed (e.g., Bellovary2025; Graham2025). These high-redshift objects may have much in common with NSCs, ultra-compact dwarfs and stripped nuclei in the local universe. The old, metal-poor NSC in M 74, with very different stellar populations from the host galaxy, may have formed either by the later migration and merging of star clusters whose stars formed at very early stages, or by very early in-situ formation. Either a combination of both, several cluster accretion episodes, or different bursts of in-situ star formation might have led to the different peaks in the SFH in Fig. 6.
@@ 7 Conclusions
We have analyzed the stellar populations of the NSC in M 74, and compared it with the host galaxy. We have performed a separate analysis, in a PSF aperture, of the NSC and the host galaxy, using a spectro-photometric decomposition method, allowing us to compare them. Additionally, we have mapped the spatially resolved stellar-population properties of the central ${\sim}3$ kpc of the host galaxy. We have reported light- and mass-weighted ages, metallicities and [Mg/Fe] abundances, and reconstructed SFHs.
@@ Appendix A Analysis of the original MUSE data cube of the galaxy
We present here the results from the analysis of the original cropped data cube of the central arcmin of M 74 (Sect. 4.1.1). The analysis was performed following the method described in Sect. 4.2. While the goal of the paper is to analyze separately the NSC and the host galaxy, the results for the integrated data cube provide a general context and serve as a reference. In Fig. 8, we show the maps of the mean age, total metallicity ([M/H]) and [Mg/Fe] abundance, from top to bottom, extracted from the original MUSE data cube. A Voronoi binning to a target S/N=100=100 was also used in this case. However, the different preprocessing compared to the host-galaxy data cube resulting from the C2D decomposition (Sect. 4.1.2) leads to a different bin size. Similarly to Fig. 5, the age map (top panel of Fig. 8) shows alternatively regions dominated by relatively young (younger than ${\sim}4$ Gyr) and old stars (older than ${\sim}8$ Gyr), roughly following the pattern of the spiral arms. [M/H] shows regions of higher and lower values. As in Fig. 5, there are regions with low metallicity in the outskirts of the mapped region, while we find supersolar metallicities within the central ${\sim}500$ pc. [Mg/Fe] abundances are in general anti-correlated with [M/H], with values relatively close to solar in the central ${\sim}500$ pc. Zooming into the very center of the middle panel in Fig. 8, a clear [M/H] drop is seen in the PSF size. However, no significant age and [Mg/Fe] differences are seen in the very center of the top and bottom panels with this spatial binning level.

### v2

@@ 6.3 NSC formation and the host-galaxy evolution
Different NSC properties have been found in galaxies of different masses (Sect. 1). This was explained with a transition in the dominant mechanisms driving NSC formation and growth, with in-situ star formation being dominant in higher mass galaxies, and globular cluster migration to the center of a galaxy dominating at lower masses (e.g., Fahrion et al. 2021, 2022b, 2022a). Among others, one piece of evidence is the transition in the NSC metallicity regime at galaxy stellar masses of about $10^{9}$ M(e.g, Neumayer et al. 2020). In more massive galaxies, NSCs tend to be more metal-rich, with [Fe/H] metallicities about $\pm 0.5$ dex around solar. For lower galaxy masses, on the other hand, NSCs have metallicities in a wide range, including very low values. In Fig.7, we have reconstructed this trend as shown in Fig. 9 of Neumayer et al. 2020, and updated it with additional measurements (NSCs from Fahrion et al. 2021, 2022a; Kacharov et al. 2018). NSCs in LTGs are shown in blue, and show, on average, lower metallicities than NSCs in ETGs (red symbols). Note also that blue triangles are total metallicities ([M/H], from Fahrion et al. 2022a), and they have to be considered upper limits for [Fe/H] metallicities. While more points will be needed to shed more light on this trend, the available points suggest that the trend of more metal-rich NSCs being hosted by more massive galaxies, clearly shown for ETGs, also holds for LTGs (but with a lower average metallicity). For higher-mass spirals, we only have four NSCs from the literature (Kacharov et al. 2018, blue circles), which appear to display lower metallicities than, on average, NSCs in ETGs of similar masses.
@@
For M 74, [Fe/H] was calculated from [M/H] and [Mg/Fe] in Table 1 following the relation ${\rm[Fe/H]}={\rm[M/H]}-0.75{\rm[Mg/Fe]}$ from Vazdekis et al. (2015). M 74’s NSC (blue open circle), is more metal poor than the rest of NSCs in Fig. 7 for similar galaxy masses, even if compared with the few LTGs. M 74 is located, in Fig. 7, in a mass regime where NSCs are thought to be dominated by in-situ formation, as suggested by their usually higher metallicities than their host galaxies (Fahrion et al. 2022b). However, its host galaxy (blue open square) is much more metal-rich than the NSC (about 0.7 dex difference in [M/H] within the central PSF region). In general, NSC old ages and low metallicities, in particular if older and more metal-poor than the host galaxy, have been interpreted as a main formation through globular cluster inspiral (e.g., Fahrion et al. 2021). If compared with other star clusters in M 74, its NSC shows similar ages to the most massive globular clusters (Whitmore et al. 2023), supporting the cluster migration scenario. On the other hand, models predict that NSCs of the mass of M 74’s NSC (${\sim}10^{7}$ M) assembled slightly more than half of their mass via in-situ star formation and the rest through globular cluster inspiral (Fahrion et al. 2022b). Recent JWST studies at high redshift have recently proposed that similar very dense objects existed already in the early universe and may have been the first structure that a galaxy formed (e.g., Bellovary 2025; Graham et al. 2025). These high-redshift objects may have much in common with NSCs, ultra-compact dwarfs and stripped nuclei in the local universe. The old, metal-poor NSC in M 74, with very different stellar populations from the host galaxy, may have formed either by the later migration and merging of star clusters whose stars formed at very early stages, or by very early in-situ formation. Either a combination of both, several cluster accretion episodes, or different bursts of in-situ star formation might have led to the different peaks in the SFH in Fig. 5.
@@ 7 Conclusions
We have analyzed the stellar populations of the NSC in M 74, and compared it with the host galaxy. We have used a spectro-photometric decomposition method, C2D, to separate the NSC light from the host galaxy. This method applies a morphological decomposition for different wavelengths, preserving variations both in the spatial and spectral directions. We performed a separate analysis, in a PSF aperture, of the NSC and the host galaxy, allowing us to compare them. Additionally, we have mapped the spatially resolved stellar-population properties of the central ${\sim}3$ kpc of the host galaxy. We have reported light- and mass-weighted ages, metallicities and [Mg/Fe] abundances, and reconstructed SFHs.
@@ Appendix A Analysis of the original MUSE data cube of the galaxy
We present here the results from the analysis of the original cropped data cube of the central arcmin of M 74 (Sect. 4.1.1). The analysis was performed following the method described in Sect. 4.2. While the goal of the paper is to analyze separately the NSC and the host galaxy, the results for the integrated data cube provide a general context and serve as a reference. In Fig. 8, we show the maps of the mean age, total metallicity ([M/H]) and [Mg/Fe] abundance, from top to bottom, extracted from the original MUSE data cube. This cube is the kinematic-corrected data cube resulting from the preparation described in Sect. 4.1.1, before being decomposed in different data cubes with C2D. A Voronoi binning to a target S/N=100=100 was also used in this case. However, the different preprocessing compared to the host-galaxy data cube resulting from the C2D decomposition (Sect. 4.1.2) leads to a different bin size. Considering that Voronoi bins are, in general, larger in Fig. 8 than in Fig. 6, the maps in the two figures are totally compatible with each other. Similarly to Fig. 6, the age map (top panel of Fig. 8) shows alternatively regions dominated by relatively young (younger than ${\sim}4$ Gyr) and old stars (older than ${\sim}8$ Gyr), roughly following the pattern of the spiral arms. [M/H] shows regions of higher and lower values. As in Fig. 6, there are regions with low metallicity in the outskirts of the mapped region, while we find supersolar metallicities within the central ${\sim}500$ pc. [Mg/Fe] abundances are in general anti-correlated with [M/H], with values relatively close to solar in the central ${\sim}500$ pc. Zooming into the very center of the middle panel in Fig. 8, a clear [M/H] drop is seen in the PSF size. However, no significant age and [Mg/Fe] differences are seen in the very center of the top and bottom panels with this spatial binning level.